Suppose that Guru and Premier produce only beats and rhymes. Guru can produce at most 400 beats or 1000 rhymes in a year. Premier can produce at most 100 beats or 200 rhymes in a year. Use the principle of comparative advantage (and maybe PPFs for the two musicians) to determine who should specialize in which activity.
Answer #1

Guru's opportunity cost of producing beats = 1000/400= 2.5 rhymes

Premier's opportunity cost of producing beats = 200/100= 2 rhymes

Guru specializes in producing rhymes and Premier specializes in producing beats